# skipabove=1em,skipbelow=1em

### v3

@@ 5.2. Feedback and Intervention
To systematically present the design of the feedback and intervention process, we do not assess whether these mechanisms rely on dedicated sensors. Rather, since all systems included in our review engage in sensor-based or perception-driven processing at some point in the ingestion-related loop, we categorize their feedback and intervention strategies in terms of modality, focusing on how information is conveyed and perceived. These modalities are classified based on the primary human sense they target, as summarized in Table 2.
Table 2. Descriptions and category of modality terms in feedback/intervention process
Modality Submodality Description Paper count
[…]
@@ 5.2.1. Visual
Visual feedback is one of the most commonly used modalities in ingestive behavior interventions, providing intuitive, accessible, and scalable means of conveying prompts, progress, or state awareness. Based on representation format, we categorize visual feedback into three submodalities: textual, graphical, and symbolic. This layered approach to visual design allows systems to balance clarity, immediacy, and ambient subtlety, tailoring feedback complexity to user needs and situational demands. A complete list of visual feedback systems and their classified modalities is available in Table 8.
Textual: Textual submodality presents information explicitly through written content such as notifications, reminders, or dietary instructions. These systems often use text-based prompts to communicate recommended actions, reflect behavioral progress, or reinforce decision-making. Examples include recommender agents (Zenun Franco, 2017; Arefeen et al., 2022), just-in-time eating assistants (Prasetyo et al., 2020; Carroll et al., 2013), and mobile self-monitoring tools (Burke et al., 2011; Krishna et al., 2025), all of which deliver messages through on-screen text. Many additional systems also employ message notifications or app-based dashboards to present dietary data or track goals in textual form, reinforcing the prominence of this submodality in behavior change design (Faltaous et al., 2021; Lee et al., 2019; Haerens et al., 2007).
Graphical: Graphical methods employ structured visual elements—such as charts, progress bars, or data visualizations—to represent behavior patterns or intervention outcomes. For instance, eat2pic translates bite patterns into abstract visual paintings as a playful form of real-time feedback (Nakamura et al., 2023), while many other systems use app interfaces to graph nutritional balance, meal composition, or eating speed, enhancing user self-awareness and reflection (Faltaous et al., 2021; Scisco et al., 2011; Lo et al., 2007).
Symbolic: Symbolic approaches convey prompts using abstract or simplified visual representations, such as color changes (Bobin et al., 2018), flashing lights (Matsui et al., 2018; Lessel et al., 2016), or emoji-style icons (Kim and Bae, 2018a), which communicate behavioral cues with minimal cognitive load. These signals are often embedded in utensils or ambient displays to indicate states like “ready to eat” or “slow down” without requiring linguistic interpretation (Agarwal et al., 2024; Kim et al., 2016b).
@@ 5.2.2. Auditory
Auditory modalities primarily rely on simple sound output devices, most commonly speakers embedded in utensils, tableware, or ambient systems. These systems use auditory feedback to deliver symbolic or affective cues that influence perception, pacing, or user awareness during meals. For instance, enhanced chewing sounds or filtered auditory effects have been applied to modulate food perception and reduce eating speed through cross-modal interaction (Kleinberger et al., 2023). Interactive forks or tableware generate rhythmic or instructive sounds, such as audio riddles or feedback tones, to encourage mindful chewing or promote healthy eating habits in children (Kadomura et al., 2013b; Cai et al., 2024). Auditory reminders are also embedded in game-based or robot-assisted scenarios to regulate intake timing (Xie et al., 2023; Kadomura et al., 2014). A complete list of studies employing auditory approaches can be found in Table 8. Compared to the visual channel, auditory cues provide a lightweight, ambient feedback pathway that can be unobtrusively integrated into multisensory interventions.
@@ 5.2.3. Haptic
Haptic modality involves bodily interventions using physical signals such as vibration or shape deformation. Instead of relying solely on visual or auditory channels, this approach engages the user’s sensorimotor experience by delivering tactile cues, modifying utensil mechanics, or altering physical resistance to shape real-time food-related actions.
Vibration: Most systems employ vibration motors or piezoelectric actuators. For instance, smart bottles vibrate when hydration targets are unmet (Faltaous et al., 2021; Varun et al., 2023), chopsticks trigger vibrations upon improper placement (Wang and Zhang, 2021), and wristbands or smartwatches vibrate in response to fast eating or insufficient chewing (Kim et al., 2016b, a; Kim and Bae, 2018b; Kamachi et al., 2023). Some systems provide vibration cues for journaling reminders (Ye et al., 2016), while others use mobile phone vibrations as hydration prompts (Chiu et al., 2009). In utensil-based designs, vibrotactile feedback is provided when eating speed exceeds a threshold (Hermans et al., 2017). Overall, vibration offers an embodied, non-disruptive alternative to visual or auditory alerts.
Deformation: Deformation-based feedback or intervention modifies the physical properties of utensils or wearable devices to guide ingestive behavior. These systems typically implement adjustable deformation through pneumatic or motorized actuation to enable conscious or subconscious intervention during eating. For example, inflatable wearables like the FatBelt simulate abdominal fullness after caloric thresholds to promote self-awareness (Pels et al., 2014), while utensil designs such as the SSpoon (Chen et al., 2022), SWAN (Khot et al., 2020), a pneumatically bending spoon (Zhang et al., 2019), or a motor-stabilized spoon for Parkinson’s patients (Singhal et al., 2024) deform or adjust posture to hinder rapid intake or support stable consumption. In contrast, other approaches rely on materials with inherent deformability, such as VR rubber donuts (Li and Bailenson, 2018) or rubbing scent-releasing tactile surfaces (Cai et al., 2024).
@@ 5.2.4. Olfactory
Olfactory modality utilizes scent stimuli to influence perception, modulate appetite, or reinforce associations with food-related behaviors. All studies related to this modality adopt the strategy of intentionally delivering odor cues to induce behavioral or perceptual changes. Methods range from digitally coordinated aroma delivery to embedded scent-triggering materials. Li and Bailenson (2018) used aromatic oil diffusers synchronized with VR scenes to emit food-related smells, enhancing satiation and reducing intake. Narumi et al. (2011) developed an olfactory display that filtered or overlaid real food odors to shift perception and promote healthier choices. Cai et al. (2024) embedded scent-infused ink in puzzle pieces that released odors through friction to support children’s food recognition. Collectively, these interventions illustrate the potential of olfactory feedback to modulate user experience on both affective and cognitive levels within ingestion-related contexts. Although relatively underexplored, scent-based strategies offer a promising channel for multisensory engagement, particularly when integrated with complementary modalities such as visual or tactile input.
@@ 5.2.5. Physiological
Among physiological modalities, electrodermal stimulation is the only technique applied in ingestion-related interventions. These systems use intraoral electrical pulses to modulate muscle activity in a closed-loop manner, responding to physiological events. Zheng et al. (2025) delivered synchronized stimulation to the mandibular vestibule based on mastication force, enhancing tactile feedback and chewing strength. Biyani et al. (2019) used tongue pressure to trigger soft palate stimulation during swallowing, aiming to reduce pharyngeal transit time and improve coordination. These sensorimotor interventions directly target neuromuscular regulation rather than perceptual awareness. Although rarely explored, such systems offer a promising path for low-latency modulation of ingestion behavior.
@@ 5.3. Design Space Mapping through Modality to Behavior
To further characterize the distribution and coverage of closed-loop approaches for sensing and intervention in ingestion health, we conducted a morphological analysis combining the dimensions of sensing and feedback/intervention modalities with their corresponding target behaviors, inspired by Zwicky (1967) and Jansen et al. (2022).
@@
This results in two-dimensional design matrices, shown in Table 3 and Table 4, which map the occurrence of each modality (x-axis) against specific behavioral targets (y-axis). To reduce redundancy and enhance interpretability, we adopt a mixed-level classification along the modality axis: for modalities with numerous fine-grained subcategories and overlapping device implementations, we collapse them into higher-level groupings to better reflect functional similarity and design relevance. The matrices enable us to identify well-covered modality–behavior pairs, underexplored combinations, and potential design opportunities. We hope this visualization can serve as a generative framework to inspire new behavior-aware intervention designs in ingestion health.
Analysis of Table 3 reveals several key trends in the current design landscape in sensing process. Human-based motion sensing, particularly wrist and object trajectory tracking, demonstrates broad behavioral coverage in oral and finger motor activities and some other context factors. Visual and acoustic modalities are also well represented, especially for food recognition, intake timing, and ingestion pace, often leveraging image, sound, or AR-based inputs.
By contrast, environment-based modalities like electrical or magnetic sensing remain underutilized, appearing in only a few cases. Physiological sensing (e.g., EMG, EDA, CGM) offers rich internal state signals but is mostly limited to specific domains such as emotional state or metabolic monitoring. This suggests potential in integrating these modalities for broader behavioral targets.
The matrix also highlights underexplored areas, including food hygiene, emotional triggers, and environmental dynamics. These gaps point to opportunities for multimodal fusion and novel device placements. Overall, the mapping reveals both saturated modality–behavior pairs and areas where new sensing strategies can enhance behavior-aware intervention design.
Sensing Modality
Human-based Environment-based
Physiological Electrical
Visual Auditory Haptic Motion Cardiac Metabolic Autonomic Visual Acoustic Force Spectral Conductivity Magnetic
Food Choices facial expression (Robinson et al., 2020; Triantafyllidis et al., 2019), gaze (VR) (Karkar et al., 2018) speech (Prasetyo et al., 2020; Arefeen et al., 2022; Luo et al., 2021; Robinson et al., 2020), intake sound (Bi et al., 2016; Maramis et al., 2014; Kalantarian et al., 2014) touch (Cai et al., 2024; Robinson et al., 2020), oral behavior (Abisha and Rajalakshmy, 2017), body weight (Harous et al., 2017; Ludwig et al., 2011; Bastida et al., 2023; Nedungadi et al., 2018) wrist trajectory (Faltaous et al., 2021; Dragoni et al., 2017; Maramis et al., 2014), Object trajectory (Kadomura et al., 2013b; Nakamura et al., 2023; Joi et al., 2016; Kadomura et al., 2014, 2013a; Bedri et al., 2022), exercise (Kumbhare et al., 2023; Mishra et al., 2024; Harous et al., 2017; Dragoni et al., 2017; Maramis et al., 2014; Bastida et al., 2023; Chukwu, 2011; Badawi et al., 2012a; Fuchs et al., 2019; Tirasirichai et al., 2018; Jung et al., 2007; Vazquez et al., 2015) blood pressure (Huang et al., 2018; Jung et al., 2007; Ludwig et al., 2011; Park and Min, 2007; Nedungadi et al., 2018), heart rate (ECG) (Jung et al., 2007; Park and Min, 2007; Badawi et al., 2012b; Nedungadi et al., 2018), blood oxygen (Park and Min, 2007; Nedungadi et al., 2018) CGM (Liang, 2024; Arefeen et al., 2024; Nedungadi et al., 2018), BIA (Balbin et al., 2020; Chukwu, 2011), cholesterol (Harous et al., 2017) temperature (Park and Min, 2007; Nedungadi et al., 2018; Jung et al., 2007) food picture (Faltaous et al., 2021; Nakamura et al., 2023; Cai et al., 2024; Kadomura et al., 2014, 2013a; Ganesh et al., 2014; Arefeen et al., 2022; YR et al., 2024; Mishra et al., 2024; Alloghani et al., 2016; Hsiao et al., 2011, 2011; Achilleos et al., 2017; Kanjalkar et al., 2024; Kwon et al., 2016; Eigen et al., 2018; Cordeiro et al., 2015; Waltner et al., 2015; Epstein et al., 2016; Akter et al., 2021), Nutrition Label (Hsiao et al., 2011; Fernandez et al., 2013; Gutiérrez et al., 2018; Fuchs et al., 2019; Volkova et al., 2016; Eyles et al., 2017; Mankoff et al., 2002; Vazquez et al., 2015) food removal (Liu et al., 2024) Meals and plate weight (de Vries et al., 2020; Joi et al., 2016; Jung et al., 2017; Zhao et al., 2021; Ritschel et al., 2018; Krishna et al., 2025; Maramis et al., 2014; Kim et al., 2016c) alcohol (Matsui et al., 2018), temperature (Bastida et al., 2023; Kwon et al., 2016; Badawi et al., 2012b), object state (Ritschel et al., 2018), food distance (Bedri et al., 2022) humidity (Bastida et al., 2023; Kwon et al., 2016; Badawi et al., 2012b)
Intake Timing intake sound (Bi et al., 2016) button click (Fortmann et al., 2014) wrist gesture (Ye et al., 2016) CGM (Liang, 2024) heart rate (Varun et al., 2023) liquid level (Varun et al., 2023) food weight (Lessel et al., 2016) temperature (Varun et al., 2023) humidity (Varun et al., 2023)
Environment indoor position (Faltaous et al., 2021)
Context Factors Activity Around Ingestion postprandial exercise (Alexander et al., 2017)
Intake Quantity body gesture (Xie et al., 2023) oral behavior (Farooq et al., 2017) bite count (Scisco et al., 2011), object trajectory (Zhou et al., 2021; Li and Bailenson, 2018) food picture (Chiu et al., 2009) leftovers (Faltaous et al., 2021), liquid level (Varun et al., 2023; Tommy et al., 2017; Ravindran et al., 2022) liquid level (Lessel et al., 2016) liquid level (Poddar et al., 2024) liquid level (Kaner et al., 2018; Kreutzer et al., 2015; Ko et al., 2007; Pankajavalli et al., 2017; Wijanarko et al., 2019), alcohol (Bobin et al., 2018) liquid level (Yıldız and Coşkun, 2019)
Ingestion Pace and Oral Motor Behavior facial expression (Kim and Bae, 2018a) intake sound (Kleinberger et al., 2023; Koizumi et al., 2011; Bi et al., 2016; Maramis et al., 2014; Kamachi et al., 2023; Chen and Yen, 2022) oral behavior (Kim et al., 2016b; Biyani et al., 2019; Zheng et al., 2025) wrist trajectory (Mendi et al., 2013; Lee et al., 2019; Scisco et al., 2011; Kim et al., 2016a; Kim and Bae, 2018b; Maramis et al., 2014; Turner-McGrievy et al., 2017), object trajectory (Kadomura et al., 2013b; Nakamura et al., 2023; Biyani et al., 2019; Hermans et al., 2017; Zhang et al., 2019), skin distance (Hori et al., 2023) wrist action (EMG) (Lee et al., 2019; Kim et al., 2016b; Sugita et al., 2018) food weight (de Vries et al., 2020; Krishna et al., 2025; Lo et al., 2007; Kim et al., 2016c; Ioakimidis et al., 2009; Mitchell et al., 2015; Hermans et al., 2017)
Food Awareness or Emotional State intake sound (Kleinberger et al., 2023; Koizumi et al., 2011; Zhang and Parker, 2020), speech (Mayumi et al., 2023; Parra et al., 2023; Zhang and Parker, 2020) wrist trajectory (Wang et al., 2025), object trajectory (Khot et al., 2020) emotion state (EDA) (Carroll et al., 2013) food picture (AR) (Narumi et al., 2011) food weight (Mitchell et al., 2015) physical contact (Han and Kang, 2017)
Target Factors and Behaviors Process Behaviors Safe Eating Practices hygiene (Wang and Zhang, 2021) temperature (Singhal et al., 2024)
Analysis of Table 4 reveals key trends in feedback modality distribution. Textual and graphical visual feedback dominate, especially for food choices, intake quantity, and ingestion pace. These channels are favored for their accessibility and clarity.
Symbolic feedback appears less frequently but with growing diversity. LED indicators are used across hydration, eating pace, and safety contexts. Symbolic auditory feedback—such as chewing sounds or metaphorical cues—emerges as a lightweight and engaging method for promoting awareness. Tactile symbolic forms, like thermochromic or pneumatic responses, remain niche but offer ambient and intuitive signaling.
Auditory feedback via speakers and alarms spans nearly all behavior types, while haptic feedback, mostly motor-based, is concentrated in emotional regulation and pace control. More advanced haptics (e.g., piezoelectric, pneumatic) remain underutilized. Experimental modalities like olfactory and physiological feedback appear in a few behaviorally rich cases, offering unique multisensory pathways.
Overall, the table highlights the dominance of visual modalities but also points to symbolic and sensory alternatives as promising design directions. Gaps in environmental, emotional, and hygiene-related feedback suggest potential for broader and more creative feedback strategies in closed-loop ingestion health systems.
@@
Notably, despite the variety of sensing pathways, few systems reuse the same sensors for both perception and intervention. Most existing feedback channels rely heavily on smartphones or static screens, rather than leveraging interactive or perceptual sensors for embodied or context-aware interventions. This reflects a substantial unexploited design space, which we further discuss and explore in the following Chapter 7.
Table 4. Mapping of feedback/intervention Modalities to Target Factors and Behaviors in Ingestion Health
Feedback/Intervention Modality
Visual
Environment microphone speaker (Faltaous et al., 2021)
Context Factors Activity Around Ingestion SMS (Alexander et al., 2017)
Target Factors and Behaviors Process Behaviors Safe Eating Practices motor vibration (Wang and Zhang, 2021; Singhal et al., 2024)

### v4

@@ 5.2. Intervention
Intervention mechanisms, whether active or passive, convey information back to users to support behavioral awareness, regulation, or change. To systematically present the design of the intervention process, we do not assess whether these mechanisms rely on dedicated sensors. Rather, since all systems included in our review engage in sensor-based or perception-driven processing at some point in the ingestion-related loop, we categorize their intervention strategies in terms of modality, focusing on how information is conveyed and perceived. To provide an overview of the types of modalities employed, Table 4 summarizes a taxonomy of intervention modalities and their corresponding publications. Accordingly, the modalities at this stage are inherently based on the primary human sense they target to deliver prompts, guidance, or other affective cues, as summarized in Table 6.
@@ 5.2.1. Visual
Visual cues is one of the most commonly used modalities in ingestive behavior interventions, providing intuitive, accessible, and scalable means of conveying prompts, progress, or state awareness. Based on representation format, we categorize visual cues into three submodalities:
i.Textual: Text-based systems convey dietary information via notifications, reminders, or instructions. They often support user behavior by prompting actions, tracking progress, and aiding decisions. Examples include recommender agents (Zenun Franco, 2017; Arefeen et al., 2022), JIT assistants (Prasetyo et al., 2020; Carroll et al., 2013), and self-monitoring apps (Burke et al., 2011; Krishna et al., 2025). Many others use dashboards or push notifications to present progress and reinforce goals (Faltaous et al., 2021; Lee et al., 2019; Haerens et al., 2007).
ii.Graphical: Graphical methods use visual elements, such as charts, bars, or data views, to convey behavioral patterns and intervention outcomes. These representations support self-awareness by illustrating nutritional balance, eating pace, or meal composition. For example, eat2pic transforms bite patterns into visual art (Nakamura et al., 2023), while others visualize ingestion summaries through app dashboards (Faltaous et al., 2021; Scisco et al., 2011; Lo et al., 2007).
iii.Symbolic: Symbolic approaches convey prompts with simplified visual representations, such as color changes (Bobin et al., 2018), flashing lights (Matsui et al., 2018; Lessel et al., 2016), or emoji-style icons (Kim and Bae, 2018a), which communicate behavioral cues with minimal cognitive load. These signals are often embedded in utensils or ambient displays to indicate states like ”ready to eat” or ”slow down (Agarwal et al., 2024; Kim et al., 2016b).
@@ 5.2.2. Auditory
Auditory modalities primarily rely on simple sound output devices, most commonly speakers embedded in utensils, tableware, or ambient systems. These systems deliver symbolic or affective auditory cues that shape perception, pacing, and awareness during meals. Such cues influence eating through cross-modal effects — for example, enhancing chewing sounds can help slow intake (Kleinberger et al., 2023). Interactive forks and tableware produce feedback tones or playful audio to encourage mindful eating, particularly in children (Kadomura et al., 2013b; Cai et al., 2024). Auditory prompts are also employed in game-based or robot-assisted contexts to guide intake timing (Xie et al., 2023; Kadomura et al., 2014). This modality provides a lightweight, ambient pathway that can be unobtrusively integrated into multisensory interventions.
@@ 5.2.3. Haptic
Haptic modality involves bodily interventions using physical signals such as vibration or deformation. Rather than relying solely on visual or auditory channels, it engages sensorimotor experience by delivering tactile cues, modifying utensil mechanics, or altering physical resistance to shape real-time food-related actions.
i.Vibration: Most systems use vibration motors or piezoelectric actuators embedded in utensils, wearables, or phones, to offer a non-disruptive alternative to visual or auditory cues. Utensil-based systems include bottles that vibrate to prompt hydration (Faltaous et al., 2021; Varun et al., 2023), chopsticks for posture correction (Wang and Zhang, 2021), and cutlery for eating speed regulation (Hermans et al., 2017). Wearables like wristbands or smartwatches deliver vibrotactile intervention for fast eating or poor chewing (Kim et al., 2016b, a; Kim and Bae, 2018b; Kamachi et al., 2023). Some systems also use phone vibrations to prompt hydration (Chiu et al., 2009) or journaling (Ye et al., 2016).
ii.Deformation: Deformative methods typically modify physical properties of devices via pneumatic or motorized actuation to influence ingestive behavior. Utensil designs like SSpoon (Chen et al., 2022), SWAN (Khot et al., 2020), and other deformable spoons (Zhang et al., 2019; Singhal et al., 2024) adjust posture or resistance to slow intake or enhance control. Wearables such as FatBelt simulate abdominal fullness through inflation after caloric thresholds (Pels et al., 2014). Other systems use deformable materials, such as VR rubber donuts (Li and Bailenson, 2018) or scent-releasing tactile surfaces (Cai et al., 2024), to influence ingestive behavior.
@@ 5.2.4. Olfactory
Olfactory modality uses scent stimuli to influence perception, appetite, or food-related associations, typically through intentional odor delivery. Methods range from digital aroma delivery to embedded scent materials, including VR-synced diffusers to curb intake (Li and Bailenson, 2018), odor filters to shift food perception (Narumi et al., 2011), and scent puzzles to aid food recognition in children (Cai et al., 2024). These approaches highlight the affective and cognitive potential of olfactory cues, especially when combined with visual or tactile input.
@@ 5.2.5. Physiological
Among physiological modalities, electrodermal stimulation is the only technique used in ingestion interventions, delivering intraoral electrical pulses to modulate muscle activity. Zheng et al. (2025) stimulated the mandibular vestibule based on chewing force, while Biyani et al. (2019) triggered soft palate stimulation via tongue pressure to enhance swallowing. Though rarely explored, such sensorimotor methods target neuromuscular control and show promise for low-latency modulation of ingestion behavior.
@@ 5.3. Design Space Mapping through Modality to Behavior
To further characterize the distribution and coverage of closed-loop paradigms for sensing and intervention in ingestion health, we conducted a morphological analysis combining the dimensions of sensing and intervention modalities with their corresponding target behaviors, inspired by Zwicky (1967) and Jansen et al. (2022).
@@
This results in two-dimensional design matrices, shown in Table 7 and Table 8, mapping each modality (x-axis) to specific behavioral targets (y-axis). To reduce redundancy and improve clarity, we adopt a mixed-level classification along the modality axis, collapsing fine-grained subcategories with overlapping implementations into higher-level groupings reflecting functional similarity. The matrices enable us to identify well-covered modality–behavior pairs, underexplored combinations, and design opportunities. This visualization can serve as a framework to inspire new behavior-aware interventions in ingestion health.
Analysis of Table 7 reveals trends in current sensing designs. Human-based motion sensing, especially wrist and object trajectory tracking, demonstrates broad behavioral coverage of oral and finger motor activities . Visual and acoustic modalities are also common for food recognition, intake timing, and pacing, often using image, sound, or AR inputs. In contrast, environment-based modalities like electrical or magnetic sensing remain underutilized. Physiological sensing (e.g., EMG, EDA, CGM) provides rich internal state signals but is mainly limited to emotional or metabolic monitoring, suggesting potential for broader integration. The matrix also highlights underexplored areas, including food hygiene, emotional triggers, and environmental dynamics. These gaps point to opportunities for multimodal fusion and novel device placements. Overall, the mapping reveals both saturated modality–behavior pairs and areas where new sensing strategies can enhance behavior-aware intervention design.
Analysis of Table 8 reveals trends in intervention modality distribution. Textual and graphical visual intervention dominate, especially for food choices, intake quantity, and ingestion pace, favored for their accessibility and clarity. Symbolic intervention appears less frequently but with increasing diversity. LED indicators are used across hydration, eating pace, and safety contexts. Symbolic auditory cues, such as chewing sounds or metaphorical signals, serve as a lightweight and engaging method to promote awareness. Tactile symbolic forms, like thermochromic or pneumatic responses, remain niche but offer ambient and intuitive signaling. Auditory interventions via speakers and alarms spans nearly all behavior types, while haptic interventions, mainly motor-based, focus on emotional regulation and pace control. More advanced haptics (e.g., piezoelectric, pneumatic) remain underused. Experimental modalities like olfactory and physiological interventions appear in a few behaviorally rich cases, offering unique multisensory pathways. Overall, the table highlights the dominance of visual modalities but also suggests symbolic and sensory alternatives as promising design directions. Gaps in environmental, emotional, and hygiene-related cues indicate potential for broader and more creative intervention strategies in closed-loop ingestion health systems.
@@
Notably, despite the variety of sensing pathways, few systems reuse the same sensors for both perception and intervention. Most existing intervention channels still rely on smartphones or static screens rather than leveraging interactive or perceptual sensors for embodied or context-aware interventions. This reflects a substantial unexploited design space, which we will discuss further in Chapter 7, including specific hypotheses about effective sensing–intervention combinations generated from the design space.